Assume that the supply of apricots is QS= 2P - 10 - a, and the demand for apricots is QD=25 - P/3 + b, where a and b are exogenous supply and demand shifters.

Assume that a=0 and b=0.

Compute the equilibrium price and quantity.
Answer #1

Answer

the demand and supply curve while a=0 and b=0 is

Q=2P-10 and Q=25-P/3

equating both as the equilibrium is at Qd=Qs

2P-10=25-P/3

2P+P/3=35

(6P+P)/3=35

7P=105

P=15

Q=2*15-10=20

the price is $15 and the quantity 20 units